### v1

# Tokenization Matters! Degrading Large Language Models through Challenging Their Tokenization

Abstract
Large Language Models (LLMs) have shown remarkable capabilities in language understanding and generation. Nonetheless, it was also witnessed that LLMs tend to produce inaccurate responses to specific queries. This deficiency can be traced to the tokenization step LLMs must undergo, which is an inevitable limitation inherent to all LLMs. In fact, incorrect tokenization is the critical point that hinders LLMs in understanding the input precisely, thus leading to unsatisfactory output. To demonstrate this flaw of LLMs, we construct an adversarial dataset, named as ADT (Adversarial Dataset for Tokenizer), which draws upon the vocabularies of various open-source LLMs to challenge LLMs’ tokenization. ADT consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto. Our empirical results reveal that our ADT is highly effective on challenging the tokenization of leading LLMs, including GPT-4o, Llama-3, Qwen2.5-max and so on, thus degrading these LLMs’ capabilities. Moreover, our method of automatic data generation has been proven efficient and robust, which can be applied to any open-source LLMs. To the best of our knowledge, our study is the first to investigating LLMs’ vulnerability in terms of challenging their token segmentation, which will shed light on the subsequent research of improving LLMs’ capabilities through optimizing their tokenization process and algorithms.
1 Introduction
In the last two years, Large Language Models (LLMs) have demonstrated remarkable capabilities in many tasks of artificial intelligence (AI), including natural language generation hoffmann2022training ; nijkamp2023codegen ; zeng2023glm130b , knowledge utilization chowdhery2022palm ; izacard2022atlas , and complex reasoning wei2023chainofthought ; zhou2023leasttomost ; kojima2023large ; taylor2022galactica . Given these capabilities, LLMs have been effectively employed by various application domains, such as healthcare tang2023does ; li2023chatdoctor ; jeblick2022chatgpt , education susnjak2022chatgpt ; bordt2023chatgpt ; Malinka_2023 , law yu2022legal ; nay2023law and so on.
Nonetheless, LLMs’ disadvantages have also been witnessed, including hallucination openai2023gpt ; bang2023multitask ; lin2022truthfulqa , knowledge recency dai2022knowledge ; Kernbach_2022 , and so on. Particularly, we also observed that for some specific queries, LLMs often produce unsatisfactory responses with words that are nonsensical, as illustrated by the two instances in Figure 1. Through checking the LLM’s tokenization results for the input sentences, we found that it is misaligned with human’s correct comprehension for the sentence. Notably, our empirical studies found that this flaw not only exists in some specific LLMs, but also is a universal issue across many mainstream LLMs. We have evaluated several prominent open-source and closed-source LLMs, including Chatglm3 zeng2023glm130b , Baichuan2 yang2023baichuan , Qwen2.5-max  [^1] https://help.aliyun.com/zh/dashscope/developer-reference/model-introduction?spm=a2c4g.11186623.0.i3, and GPT-4o  [^2] https://platformopenai.com/docs/models/gpt4o. Our experiment results reveal that regardless of LLMs’ scales or their claimed capabilities, they inevitably generate incorrect or entirely nonsensical responses for some specific inputs when their tokenization results for the input sentences are obviously wrong. Consequently, we believe that LLMs’ tokenization errors prevent them from accurately understanding the input text, leading to their incorrect responses.
Figure 1: Two instances of LLM generating incorrect response due to incorrect tokenization. Case 1 is a Chinese input instance, of which the English translation is noted below according to its correct tokenization. In Case 2, a space is omitted between ‘moves’ and ‘table’, causing the LLM’s incorrect tokenization, which is detailed in Section 3.2.
As we know, LLMs’ tokenization flaws stem from the algorithms of their tokenizers, most of which are based on subword-level vocabularies. The popular tokenization algorithms include Byte-Pair Encoding (BPE) sennrich2016neural , WordPiece Schuster_Nakajima_2012 , and Unigram kudo2018subword . However, no vocabulary can perfectly cover all possible ways of various expressions in the inputs. When applying the greedy principle, the algorithms are easy to generate unsatisfactory results which are not fully aligned with the correct intention of users’ input. Unfortunately, in cases of tokenization errors, all subsequent optimization operations for LLMs cannot completely solve this underlying problems caused by their tokenization algorithms.
In the domain of natural language processing (NLP), the existing studies related to tokenization primarily focus on refining or optimizing various tokenization algorithms. Meanwhile, the discussions on LLMs’ vulnerability including attack or challenge techniques, have been more concerned with the security of LLMs. About the challenges posed by tokenization deficiencies, only Sander and Max land2024fishing have discussed this issue from the perspective of under-trained tokens in LLMs. It is worth noting that, to the best of our knowledge, there has been no research specifically focusing on the unsatisfactory token segmentation of LLMs, which is indeed a critical concern causing LLMs’ vulnerability.
In this paper, we focus on the critical flaw in LLMs’ tokenization process, and try to reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses for some specific queries. To this end, for the first time, we construct a dataset, namely ADT (Adversarial Dataset for Tokenizer), to challenge the tokenization of different representative LLMs. Our ADT dataset consists of two subsets: the manually constructed ADT-Human and automatically generated ADT-Auto. At first, we export the vocabularies from multiple mainstream LLMs, based on which ADT-Human is constructed. Our experiment results demonstrate that ADT-Human can effectively challenge the LLMs’ tokenization, leading to their completely incorrect responses. Furthermore, we also develop a framework for automatically generating adversarial data to construct the dataset more efficiently. Initially, we export an LLM’s vocabulary and identify the trap words that can influence the model’s performance. By inputting these trap words into GPT-4 openai2023gpt with prompt, we leverage its capability to get available instances, which are then inspected manually to ensure their quality. With minimal human effort, we successfully construct ADT-Auto with 231 instances, validating the effectiveness of our framework and highlighting the inevitable flaws in LLMs’ tokenization once again.
In summary, the contributions of this paper are as follows:
  1. 1.
    For the first time, we investigate LLMs’ vulnerability for some special inputs in terms of challenging their token segmentation, providing a new perspective of studying LLMs’ disadvantages.
  2. 2.
    We propose an effective framework to construct a new dataset ADT, which can be used to challenge various LLMs’ tokenization, exposing their vulnerability for specific queries.
  3. 3.
    Our experiment results obviously reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses, which can shed light on the subsequent work of improving LLMs through optimizing their tokenization.
2 Related work
Algorithms for tokenization.
Tokenization is a basic but crucial step in NLP. In order to get the models to better understand the text, great improvements have been made to the tokenization process. Currently, the mainstream tokenization approachs are based on sub-word units, including Byte Pair Encoding (BPE), WordPiece, and Unigram. BPE sennrich2016neural forms an extensive vocabulary starting from the base of character-level tokens, merging these token pairs and intercalating them into the vocabulary. The merging rule is to select the adjacent pairs with the highest word frequency. LLMs including GPT-3 brown2020language , RoBERTa liu2019roberta , BART lewis2019bart , Llama2 touvron2023llama , and Baichuan yang2023baichuan are based on BPE. WorldPiece Schuster_Nakajima_2012 is similar to BPE, but it differs in the strategy of merging pairs with reference to mutual information rather than frequency. LLMs built upon WordPiece include BERT devlin2019bert , DistilBERT sanh2020distilbert , and Electra clark2020electra . Different from these two algorithms, Unigram kudo2018subword starts with a large vocabulary and gradually trims it down to a smaller one. It measures the importance of subwords by computing the loss incurred by deleting different subwords, retaining those with high significance. LLMs utilizing Unigram include AlBERT lan2020albert and mBART liu2020multilingual . Besides, many tokenization tools integrating these algorithms have been developed, such as Google’s SentencePiece kudo2018sentencepiece and OpenAI’s tiktoken [^3] https://github.com/openai/tiktoken, which simplify the process of LLMs tokenization greatly. However, there has been no work concerning the shared risks inherent in these algorithms. So we firstly focus on this issue, discussing the underlying risks of LLMs’ tokenization.
Attack techniques in LLMs.
With the growing prominence of LLMs, the security and vulnerability of these models have attracted significant attention, and even advanced LLMs like GPT-4 are no exception. A surge of research in this field is underway, with researchers launching attacks on LLMs from various aspects esmradi2023comprehensive ; chowdhury2024breaking including but not limited to, Prompt Injection choi2022prompt , Model Theft krishna2020thieves , Data Reconstruction carlini2021extracting , Data Poisoningwallace2021concealed ; xu2024instructions , and Member Inference Attack 7958568 . For instance, Prompt Injection Attack refers to a scenario where an attacker crafts malicious prompts with the intention of deceiving a language model to generate outputs inconsistent with its training data and anticipated functionality. Associating with PI attacks, threat actors aim for information gathering, fraud, intrusion, malware, content manipulation, and availability attacks choi2022prompt . Carlini, Nicholas, et al carlini2021extracting executed the Data Reconstruction attack on GPT-2, extracting personal identity information, code, and UUIDs. Data Poisoning pertains to the deliberate introduction of corrupted or malicious data into the training dataset to manipulate the model’s behavior. Member Inference Attack indicates that models tend to memorize sensitive information, posing them as potential targets for attacks. A study postulates that an attacker can train a binary classifier using the input-output pairs from LLMs to determine whether a specific individual was part of the model’s training set 7958568 . Diverging from existing works, our research innovatively suggests attacking the capabilities of language models from the perspective of Tokenization.
3 Methodology of dataset construction
In this section, we describe the process of constructing ADT (Adversarial Dataset for Tokenizer) in detail, which is used to challenge LLMs’ tokenization and thus reveal LLMs’ vulnerability. Our ADT contains two subsets, the manually constructed ADT-Human and automatically generated ADT-Auto.
3.1 Vocabulary export
In fact, an instance in our dataset comprises one sentence containing the token (word) that could challenge LLMs’ tokenization and one question related to the sentence. To inspect whether LLMs can accurately recognize these challenging tokens within various contexts based on their memories, the tokens should come from the vocabularies of LLMs. Therefore, the first step of our dataset construction is to export the vocabularies of LLMs. Given that Chinese is more complex and challenging than English in terms of tokenization, as detailed in Section 3.2, we pay more attention to the issues in Chinese data.
Specifically, we select five widely-used open-source LLMs which are trained on extensive Chinese corpus, to export their Chinese vocabularies, including Chatglm3-6B zeng2023glm130b , Baichuan2-13B-Chat yang2023baichuan , Yi-34B-Chat ai2024yi , Qwen-7B-Chat bai2023qwen , and Qwen1.5-72B-Chat bai2023qwen . Besides, we export the English vocabularies from three English-based LLMs, i.e., Llama-3-8B-Instruct touvron2023llama , Llama-3-70B-Instruct touvron2023llama , and Mixtral-8x7B-Instruct-v0.1 jiang2024mixtral . Notably, the Chinese vocabulary of Qwen-7B-Chat and Qwen1.5-72B-Chat is the same, so is the English vocabulary of Llama-3-8B-Instruct and Llama-3-70B-Instruct. These LLMs use SentencePiece or tiktoken to train their tokenizers, which are tokenizer tools widely used by various language models.
In the process of exporting the vocabularies, a sequence of straightforward operations is considered. Initially, the tokenizer is decoded to obtain the vocabulary, followed by the removal of leading and trailing spaces from each token. Notably, if SentencePiece is used for tokenization in the training phase, some certain tokens may begin with a special token ‘_’ because SentencePiece treats the input text just as a sequence of Unicode characters, and whitespace is also handled as a normal symbol. To handle the whitespace as a basic token explicitly, SentencePiece first escapes the whitespace with a meta symbol ‘_’ (U+2581)  [^4] https://github.com/google/sentencepiece. Consequently, when exporting the vocabularies for models trained with SentencePiece, we replace ‘_’ with a blank character.
We summarize the exported vocabulary sizes of different LLMs in Table 1.
Table 1: Vocabulary sizes of different LLMs.
Language Model Vocabulary Size Vocabulary Size
of Specific Language
Chinese Chatglm3-6B 64,789 30,922
Baichuan2-13B-Chat 125,696 70,394
Yi-34B-Chat 64,000 21,353
Qwen-7B-Chat 151,851 24,953
Qwen1.5-72B-Chat 151,646 24,953
English Llama-3-8B-Instruct 128,257 72,420
Llama-3-70B-Instruct 128,257 72,420
Mixtral-8x7B-Instruct-v0.1 32,000 25,056
3.2 ADT-Human construction
Based on the exported vocabularies, we manually construct our dataset ADT-Human, to challenge and evaluate the tokenization of different LLMs. The process of manual construction is depicted in Figure 2.
Figure 2: Our framework of constructing ADT-Human manually.
Specifically, we sophisticatedly select eleven tokens from each Chinese vocabulary, and seven tokens for each English vocabulary. Then, for each selected token, we adopt one of the three approaches listed in Table 2, to convert it into a challenging span through inserting a special character span before or (and) after it. These challenging spans would disrupt the conventional tokenization process, thus confusing LLMs. The considerations in the three conversion approaches are introduced as follows.
  1. 1.
    Before: A character span $s$ is inserted before the origin token, causing that the concatenation of $s$ and the token’s prefix is just another token (word) existing in the vocabulary, as ‘move’+‘s’$\rightarrow$‘moves’ in Table 2.
  2. 2.
    After: A character span $s$ is inserted after the origin token, causing that the concatenation of the token’s suffix and $s$ is just another token existing in the vocabulary, as ‘n’+‘othing’$\rightarrow$‘nothing’ in Table 2.
  3. 3.
    Before & After: Character spans $s_{1}$ and $s_{2}$ are inserted before and after the origin token, respectively, causing that the concatenation of $s_{1}$ and the token’s prefix is just another token existing in the vocabulary, meanwhile so is the concatenation of the token’s suffix and $s_{2}$, as ‘six’+‘th’$\rightarrow$‘sixth’ and ‘is’+‘land’$\rightarrow$‘island’ in Table 2.
Next, for each challenging span $s_{c}$, we manually compose a corresponding instance as the data in ADT-Human. One instance consists of a sentence in which $s_{c}$ presents, and a corresponding question of which the answer comes from $s_{c}$. In Appendix A, we list all instances (along with their correct tokenization) in ADT-Human for the four Chinese vocabularies and two English vocabularies we export. As we can see, it is challenging for the LLMs to understand the instances due to the presence of challenging spans.
Table 2: Three approaches of generating challenging spans.
Approach Origin token After insertion Model tokenization Human tokenization
(Challenging span)
Before Schema BC ABC [A, BC] [AB, C]
Example stable movestable [move, stable] [moves, table]
After Schema AB ABC [AB, C] [A, BC]
Example driven drivenothing [driven, othing] [drive, nothing]
Before & After Schema BC ABCD [A, BC, D] [AB, CD]
Example this sixthisland [six, this, land] [sixth, island]
It’s worth noting that the tokenization difficulty of English is less than that of Chinese, since spaces are regularly used as delimiters to separate each word from others in English. Moreover, affixation is common in English word structure, enabling tokenizers to divide a single word into several sections, which can help avoid incorrect tokenization to some extent. Thus, during the manual construction of English instances, we deliberately exclude the spaces between some tokens to provoke challenges to tokenization process.
3.3 ADT-Auto generation
Given the inefficiency of constructing dataset manually, we further develop an automatic construction framework for the dataset to challenge LLMs’ tokenization.
As discussed in Section 3.2, the challenges of tokenization in English are less severe than in Chinese. Consequently, we primarily concentrate on the automatic generation of dataset in Chinese. Figure 3 illustrates the process of automatically constructing our dataset ADT-Auto.
Figure 3: Our framework of generating ADT-Auto automatically.
3.3.1 Word-pair matching
The automatic generation of dataset is also based on the exported vocabularies. From the vocabularies, we first seek some qualified word pairs. Given two words Word 1 and Word 2, they are considered to be match when meeting the following criteria: The suffix (or whole) of Word 1 can be concatenated with the prefix (or whole) of Word 2, as a token (word) existing in the vocabulary, denoted as Trap Word. Meanwhile, the concatenation of Word 1 and Word 2 should not be a token existing in the vocabulary.
Accordingly, there are also three situations related to Word 1, Word 2 and Trap Word, corresponding to three scenarios in Table 2.
  1. 1.
    If Word 1 is included by Trap Word, this situation corresponds to the schema inserting a span after Trap Word.
  2. 2.
    If Word 2 is included by Trap Word, this situation corresponds to the schema inserting a span before Trap Word.
  3. 3.
    If neither Word 1 nor Word 2 is included by Trap Word, this situation corresponds to the schema inserting a span both before and after Trap Word.
Above clarifications indicate that the concatenation of Word 1 and Word 2 (as ‘ABCD’ in Figure 3) corresponds to the challenging span in the construction process of ADT-Human. Our objective is to compose a sentence that can not only convey the semantic essence of both Word 1 and Word 2, but also induce LLM to isolate Trap Word during the tokenization process. Notably, we neglect the situation that the concatenation of Word 1 and Word 2 is just Trap Word, because it would not pose challenges to LLMs’ tokenization.
To augment the efficacy of Trap Word upon matching, we also consider the word-pair matching criterion that the remaining parts of Word 1 and Word 2 excluded by Trap Word should also exist in the vocabulary as a token (word). Furthermore, the first or last character of Trap Word cannot be a Chinese stop character.
3.3.2 Instance generalization
With each Word 1, Word 2 and their corresponding Trap Word, we harness GPT-4 to generate an instance of ADT-Auto, which consists of a sentence and a corresponding question. Specifically, we prompt GPT-4 to generate a sentence that is used to challenge the tokenization of LLMs. To this end, the generated sentence is required to include the concatenation of Word 1 and Word 2 (inevitably including Trap Word). In addition, GPT-4 is also required to devise a question related to the generated sentence, which is used to evaluate LLMs’ tokenization performance through their answers. The instance must meet the criteria as below: The sentence should convey the meanings of both Word 1 and Word 2, while the question’s answer should come from Word 1, Word 2, Trap Word or their combination. Thus, the influence of LLMs’ incorrect tokenization can be identified by checking their answers to the question.
The prompt for GPT-4 to generate ADT-Auto instances also includes some demonstration examples in addition to the task instructions, which is illustrated in Appendix B.
3.3.3 Filtering based on Trap Word
The goal of our dataset is to expose the flaw of LLMs’ tokenization. Therefore, we will only retain the instances that can induce the tokenization problem of LLMs. To this end, we check each instance generated at the previous step. For these instances, the presence of Trap Word implies a challenging case that is likely to induce tokenization problems. Accordingly, we first select an open-source LLM and get the Trap Word set based on its vocabulary. Given an instance, we retain it if a certain Trap Word is both in the LLM’s tokenization list for the instance’s sentence (as List 1 in Figure 3) and in its tokenization list for the answer (as List 2 in Figure 3). Such filtering criterion indicates that the LLM commits tokenization errors on understanding the sentence and response for the instance. Thus, this instance is also likely to challenge other LLMs’ tokenization.
3.3.4 Manual filtering
To ensure that the retained instances can induce tokenization problems and meanwhile have reasonable expressions, we further adopt the manual assessment for instances. Specifically, we further select the high-qualify instances with considering their sentence expressions and LLMs’ responses. Notably, we might still retain some instances to which the used LLM has correct response, since the other LLMs are still likely to commit inaccurate tokenization for these instances, resulting in unsatisfactory responses.
Due to space limitation, we take Qwen-7B as an example to illustrate the process of generating ADT-Auto instances. There are 24,953 Chinese tokens in Qwen-7B’s vocabulary, and after word-pair matching, 1,764,692 word-pairs are obtained. From the matching word-pairs, 8,000 pairs are selected randomly and then used for instance generation by GPT-4. Due to the inherent stochastic characteristic of LLMs on response generation, we conduct three iterations of filtering operations introduced in Subsection 3.3.3 to get more qualified instances, and thus retain 894 instances. Next, through the manual filtering introduced in Subsection 3.3.4, we retain 231 for ADT-Auto in the end.
4 Experiments
4.1 Experiment setup
Considering the open-source LLMs used in our experiments, we select the LLMs previously used in the construction of ADT-Human, including Chatglm3-6B, Baichuan2-13B-Chat, Yi-34B-Chat, Qwen-7B-Chat, and Qwen1.5-72B-Chat for Chinese data, as well as Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1 for English data. We test these LLMs using both locally deployed versions and API versions. For the closed-source LLMs used in our experiments, we test GPT-4o, GPT-4, GPT-3.5-Turbo, Qwen2.5-max, step-1-8k [^5] https://platform.stepfun.com/docs/llm/text, moonshot-v1-8k [^6] https://platform.moonshot.cn/docs, ERNIE-3.5-8K777 https://cloud.baidu.com/doc/WENXINWORKSHOP/s/jlil56u11 for Chinese data, and GPT-4o, GPT-4 and GPT-3.5-Turbo for English data.
In our experiments, we directly use the dataset ADT constructed with the method introduced in Section 3, which includes the manually constructed ADT-Human (containing Chinese and English instances) and the automatically generated ADT-Auto (only containing Chinese instances).
In addition, we conduct our experiments on the platform with four A800 GPUs.
4.2 ADT-Human’s challenges to LLMs
Firstly, we investigate the challenges posed by our manually constructed dataset ADT-Human to LLMs. To this end, we evaluate the experimented LLMs’ performance on the dataset, through counting the number of incorrect answers generated by the LLMs for the questions in the instances. Recalling the process of manually composing challenging spans introduced in Section 3.2, the span ‘BC’ in Table 2 is in fact a Trap Word mentioned in Section 3.3. Hence, for a given LLM, its answer including a Trap Word is identified as its inaccurate response for the question undoubtedly. Such a case also indicates that the LLM generates a wrong token segmentation.
The percentage of incorrect responses provided by the LLMs for the Chinese data (corresponding to the four LLMs’ vocabularies) and English data (corresponding to the two LLMs’ vocabularies) are presented in Table 3 and Table 4, respectively. It should be noted that there is no API version of Chatglm3-6B. The results shows that the ADT-Human poses a significant challenge to both open-source and closed-source LLMs on their tokenization, resulting in very high rates of inaccurate responses.
Table 3: Error rates(%) of answers on ADT-Human (Chinese).
Model Source LLM of vocabulary Overall error rate Chatglm3 Baichuan2 Yi Qwen Open-source (Local) Chatglm3-6B 100.00 100.00 100.00 90.91 97.73 Baichuan2-13B-Chat 90.91 100.00 81.82 100.00 93.18 Yi-34B-Chat 72.73 63.64 100.00 100.00 84.09 Qwen-7B-Chat 100.00 72.73 90.91 100.00 90.91 Qwen1.5-72B-Chat 90.91 45.45 81.82 100.00 79.55 Open-source (API) Baichuan2-13B-Chat 100.00 100.00 90.91 100.00 97.73 Yi-34B-Chat 81.82 54.55 100.00 90.91 81.82 Qwen-7B-Chat 100.00 81.82 81.82 100.00 90.91 Qwen1.5-72B-Chat 90.91 54.55 81.82 100.00 81.82 Closed-source GPT-4o 72.73 27.27 54.55 90.91 61.36 GPT-4 81.82 45.45 45.45 27.28 50.00 GPT-3.5-Turbo 72.73 27.27 72.73 72.73 61.36 Qwen2.5-max 90.91 72.73 90.91 100.00 88.64 step-1-8k 63.64 18.18 72.73 63.64 54.55 moonshot-v1-8k 81.82 27.27 100.00 81.82 72.73 ERNIE-3.5-8K 72.73 54.55 54.55 72.73 63.64
Table 4: Error rates(%) of answers on ADT-Human (English).
Model Source LLM of vocabulary Overall error rate Llama-3 Mixtral Open-source (Local) Llama-3-8B-Instruct 100.00 85.71 92.86 Llama-3-70B-Instruct 57.14 71.43 64.29 Mixtral-8x7B-Instruct-v0.1 85.71 100.00 92.86 Open-source (API) Llama-3-8B-Instruct 100.00 71.43 85.71 Llama-3-70B-Instruct 57.14 28.57 42.86 Mixtral-8x7B-Instruct-v0.1 71.43 100.00 85.71 Closed-source GPT-4o 57.14 71.43 64.29 GPT-4 57.14 57.14 57.14 GPT-3.5-Turbo 71.43 57.14 64.29
It is also worth mentioning that the recent state-of-the-art (SOTA) GPT-4o cannot outperform GPT-4 (higher error rate), implying that the advancements of these LLMs have not yet addressed this primary but challenging problem.
4.3 ADT-Auto’s challenges to LLMs
Similar to the investigation of ADT-Human’s challenges to LLMs’ tokenization, we also evaluate the LLMs’ performance on ADT-Auto. We still test five open-source LLMs with their both locally deployed and API versions, and seven closed-source LLMs only having APIs.
ADT-Auto’s instances come from Qwen-7B’s vocabulary, and the rates of LLMs’ inaccurate responses for these instances’ questions are listed in Table 5. Based on the results, we have the following observations and analysis:
Table 5: Error rates of answers on ADT-Auto.
Model Fraction Error rate(%)
Open-source (Local) Chatglm3-6B 156/231 67.53
Baichuan2-13B-Chat 147/231 63.64
Yi-34B-Chat 90/231 38.96
Qwen-7B-Chat 185/231 80.09
Qwen1.5-72B-Chat 93/231 40.26
Open-source (API) Baichuan2-13B-Chat 167/231 72.29
Yi-34B-Chat 80/231 34.63
Qwen-7B-Chat 160/231 69.26
Qwen1.5-72B-Chat 97/231 41.99
Closed-source GPT-4o 75/231 32.47
GPT-4 89/231 38.53
GPT-3.5-Turbo 99/231 42.86
Qwen2.5-max 75/231 32.47
step-1-8k 60/231 25.97
moonshot-v1-8k 60/231 25.97
ERNIE-3.5-8K 51/231 22.08
  1. 1.
    Compared with the closed-source LLMs, the open-source LLMs suffer from ADT-Auto’s challenges more apparently. It implies that these closed-source LLMs, as the profit-making flagships of their creator companies, naturally have stronger capabilities than the open-source LLMs that are created for public usage.
  2. 2.
    Compared with ADT-Human, ADT-Auto is less challenging to LLMs, since the sentences generated by GPT-4 have more formal, regular or simple syntaxes and expressions than the manually composed sentences in ADT-Human. Thus, these sentences in ADT-Auto are relatively easy for LLMs’ understanding.
  3. 3.
    Qwen1.5-72B-Chat has much lower error rates than Qwen-7B-Chat, although they have the same vocabulary. We specially check the instances to which Qwen1.5-72B-Chat gives the correct answers but Qwen-7B-Chat gives the wrong answers, and find the two models have the same incorrect tokenization lists for these instances. It suggests that their different performance is not caused by their tokenization. These results also imply that in the case of incorrect tokenization, the bigger LLMs are more robust and likely to generate correct responses than the smaller LLMs thanks to their stronger capabilities brought by the larger scale.
5 Perspectives and conclusion
Based on our findings in this paper, for the work dedicating to improve LLMs’ tokenization, and so as to improve LLMs’ performance on varied tasks, it is reasonable and promising to expand LLMs’ vocabulary sizes or developing more powerful tokenizers or effective tokenization algorithms. As the idea proposed by Kudo kudo2018subword , employing multiple subword segmentation methods might be a feasible strategy to enhance the robustness of LLMs’ tokenization process. We hope that our work in this paper and the dataset could shed light on the subsequent research on improving LLMs’ tokenization and then the performance on other tasks.
In this paper, we dedicate to deeply investigating the relationship between LLMs’ vulnerability on tokenization and their unsatisfactory responses for some tasks. To this end, we construct an adversarial dataset ADT (Adversarial Dataset for Tokenizer) consisting of a manual constructed subset ADT-Human and an automatically generated subset ADT-Auto, of which each instance includes a sentence and a corresponding question. Our experiments obviously demonstrate that our dataset does challenge the studied open-source and closed-source LLMs’ token segmentation, resulting in their incorrect answers to the instances’ questions.
Appendix A Details of ADT-Human
A.1 Chatglm3
A.2 Baichuan2
A.3 Yi
A.4 Qwen
A.5 Llama-3
A.6 Mixtral
Appendix B Details of prompt

### v2

@@
 In fact, incorrect tokenization is the critical point that hinders LLMs in understanding the input precisely, thus leading to unsatisfactory output.
+This defect is more obvious in Chinese scenarios.
 To demonstrate this flaw of LLMs, we construct an adversarial dataset, named as ADT (Adversarial Dataset for Tokenizer), which draws upon the vocabularies of various open-source LLMs to challenge LLMs’ tokenization.
 ADT consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto.
-Our empirical results reveal that our ADT is highly effective on challenging the tokenization of leading LLMs, including GPT-4o, Llama-3, Qwen2.5-max and so on, thus degrading these LLMs’ capabilities.
+Our empirical results reveal that our ADT is highly effective on challenging the tokenization of leading LLMs, including GPT-4o, Llama-3, Deepseek-R1 and so on, thus degrading these LLMs’ capabilities.
 Moreover, our method of automatic data generation has been proven efficient and robust, which can be applied to any open-source LLMs.
-To the best of our knowledge, our study is the first to investigating LLMs’ vulnerability in terms of challenging their token segmentation, which will shed light on the subsequent research of improving LLMs’ capabilities through optimizing their tokenization process and algorithms.
+In this paper, we substantially investigate LLMs’ vulnerability in terms of challenging their token segmentation, which will shed light on the subsequent research of improving LLMs’ capabilities through optimizing their tokenization process and algorithms.
+Tokenization Matters! Degrading Large Language Models through Challenging Their Tokenization
+Dixuan Wang1, Yanda Li1, Junyuan Jiang2, Zepeng Ding1, Ziqin Luo1, Guochao Jiang1, Jiaqing Liang1thanks: Corresponding Authors, Deqing Yang [^1] footnotemark: 1 School of Data Science, Fudan University, Shanghai, China 1 School of Management, Fudan University, Shanghai, China 2 {dxwang23,ydli22,jiangjy21,zqluo22,gcjiang22}@m.fudan.edu.cn, {dingzepeng,liangjiaqing,yangdeqing}@fudan.edu.cn
 1 Introduction
-In the last two years, Large Language Models (LLMs) have demonstrated remarkable capabilities in many tasks of artificial intelligence (AI), including natural language generation hoffmann2022training ; nijkamp2023codegen ; zeng2023glm130b , knowledge utilization chowdhery2022palm ; izacard2022atlas , and complex reasoning wei2023chainofthought ; zhou2023leasttomost ; kojima2023large ; taylor2022galactica .
-Given these capabilities, LLMs have been effectively employed by various application domains, such as healthcare tang2023does ; li2023chatdoctor ; jeblick2022chatgpt , education susnjak2022chatgpt ; bordt2023chatgpt ; Malinka_2023 , law yu2022legal ; nay2023law and so on.
-Nonetheless, LLMs’ disadvantages have also been witnessed, including hallucination openai2023gpt ; bang2023multitask ; lin2022truthfulqa , knowledge recency dai2022knowledge ; Kernbach_2022 , and so on.
-Particularly, we also observed that for some specific queries, LLMs often produce unsatisfactory responses with words that are nonsensical, as illustrated by the two instances in Figure 1.
-Through checking the LLM’s tokenization results for the input sentences, we found that it is misaligned with human’s correct comprehension for the sentence.
+In the last two years, Large Language Models (LLMs) have demonstrated remarkable capabilities in many tasks of artificial intelligence (AI), including natural language generation Hoffmann et al.
+(2022); Nijkamp et al.
+(2023); Zeng et al.
+(2023), knowledge utilization Chowdhery et al.
+(2023); Izacard et al.
+(2023), and complex reasoning Wei et al.
+(2022); Zhou et al.
+(2023); Kojima et al.
+(2022); Taylor et al.
+(2022).
+Given these capabilities, LLMs have been effectively employed by various application domains, such as healthcare Tang et al.
+(2023); Li et al.
+(2023); Jeblick et al.
+(2022), education Susnjak (2022); Bordt and von Luxburg (2023); Malinka et al.
+(2023), law Yu et al.
+(2022); Nay (2022) and so on.
+Figure 1: Two instances of LLM generating incorrect response due to incorrect tokenization.
+Case 1 is a Chinese input instance, of which the English translation is noted below according to its correct tokenization.
+In Case 2, a space is omitted between ‘sing’ and ‘lemon’, causing the LLM’s incorrect tokenization, which is detailed in Section 3.2.
+Nonetheless, LLMs’ disadvantages have also been witnessed, including hallucination OpenAI (2023); Bang et al.
+(2023); Lin et al.
+(2022), knowledge recency Dai et al.
+(2022); Kernbach (2022), and so on.
+Particularly, we observed that for some specific queries, LLMs often produce unsatisfactory responses with words that are nonsensical, as illustrated by the two instances in Figure 1.
+Through checking the LLM’s tokenization results for input sentences, we found that it is misaligned with human’s correct comprehension for the sentences.
 Notably, our empirical studies found that this flaw not only exists in some specific LLMs, but also is a universal issue across many mainstream LLMs.
-We have evaluated several prominent open-source and closed-source LLMs, including Chatglm3 zeng2023glm130b , Baichuan2 yang2023baichuan , Qwen2.5-max  [^1] https://help.aliyun.com/zh/dashscope/developer-reference/model-introduction?spm=a2c4g.11186623.0.i3, and GPT-4o  [^2] https://platformopenai.com/docs/models/gpt4o.
+We have evaluated several prominent open-source and closed-source LLMs, including Chatglm3 Zeng et al.
+(2023), Qwen2.5-max Team (2024), Deepseek-R1 DeepSeek-AI (2025), and GPT-4o OpenAI (2024).
 Our experiment results reveal that regardless of LLMs’ scales or their claimed capabilities, they inevitably generate incorrect or entirely nonsensical responses for some specific inputs when their tokenization results for the input sentences are obviously wrong.
 Consequently, we believe that LLMs’ tokenization errors prevent them from accurately understanding the input text, leading to their incorrect responses.
-Figure 1: Two instances of LLM generating incorrect response due to incorrect tokenization.
-Case 1 is a Chinese input instance, of which the English translation is noted below according to its correct tokenization.
-In Case 2, a space is omitted between ‘moves’ and ‘table’, causing the LLM’s incorrect tokenization, which is detailed in Section 3.2.
 As we know, LLMs’ tokenization flaws stem from the algorithms of their tokenizers, most of which are based on subword-level vocabularies.
-The popular tokenization algorithms include Byte-Pair Encoding (BPE) sennrich2016neural , WordPiece Schuster_Nakajima_2012 , and Unigram kudo2018subword .
+The popular tokenization algorithms include Byte-Pair Encoding (BPE) Sennrich et al.
+(2016), WordPiece Schuster and Nakajima (2012), and Unigram Kudo (2018).
 However, no vocabulary can perfectly cover all possible ways of various expressions in the inputs.
-When applying the greedy principle, the algorithms are easy to generate unsatisfactory results which are not fully aligned with the correct intention of users’ input.
+The algorithms may potentially generate unsatisfactory results which are not fully aligned with the true intention of users’ input.
 Unfortunately, in cases of tokenization errors, all subsequent optimization operations for LLMs cannot completely solve this underlying problems caused by their tokenization algorithms.
@@
 Meanwhile, the discussions on LLMs’ vulnerability including attack or challenge techniques, have been more concerned with the security of LLMs.
-About the challenges posed by tokenization deficiencies, only Sander and Max land2024fishing have discussed this issue from the perspective of under-trained tokens in LLMs.
-It is worth noting that, to the best of our knowledge, there has been no research specifically focusing on the unsatisfactory token segmentation of LLMs, which is indeed a critical concern causing LLMs’ vulnerability.
+For LLMs, in terms of the challenges posed by tokenization deficiencies, Sander and Max Land and Bartolo (2024) have discussed this issue from the perspective of under-trained tokens in LLMs.
+It is worth noting that, to the best of our knowledge, there has been no research specifically focusing on the unsatisfactory token segmentation of LLMs, particularly in Chinese scenarios, which is indeed a critical concern causing LLMs’ vulnerability.
 In this paper, we focus on the critical flaw in LLMs’ tokenization process, and try to reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses for some specific queries.
-To this end, for the first time, we construct a dataset, namely ADT (Adversarial Dataset for Tokenizer), to challenge the tokenization of different representative LLMs.
-Our ADT dataset consists of two subsets: the manually constructed ADT-Human and automatically generated ADT-Auto.
+To this end, for the first time, we construct a dataset, namely ADT (Adversarial Dataset for Tokenizer), to challenge the tokenization of various LLMs.
+ADT dataset consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto.
 At first, we export the vocabularies from multiple mainstream LLMs, based on which ADT-Human is constructed.
-Our experiment results demonstrate that ADT-Human can effectively challenge the LLMs’ tokenization, leading to their completely incorrect responses.
-Furthermore, we also develop a framework for automatically generating adversarial data to construct the dataset more efficiently.
-Initially, we export an LLM’s vocabulary and identify the trap words that can influence the model’s performance.
-By inputting these trap words into GPT-4 openai2023gpt with prompt, we leverage its capability to get available instances, which are then inspected manually to ensure their quality.
+Our experiment results demonstrate that ADT-Human can effectively challenge LLMs’ tokenization, leading to their completely incorrect responses.
+Furthermore, we also develop a framework for automatically generating adversarial data to construct dataset more efficiently.
+Initially, we export LLM’s vocabulary and identify the trap words that can influence the model’s performance.
+By inputting these trap words into GPT-4 OpenAI (2023) with prompt, we leverage its capability to get available instances, which are then inspected manually to ensure quality.
 With minimal human effort, we successfully construct ADT-Auto with 231 instances, validating the effectiveness of our framework and highlighting the inevitable flaws in LLMs’ tokenization once again.
@@
   1. 1.
-    For the first time, we investigate LLMs’ vulnerability for some special inputs in terms of challenging their token segmentation, providing a new perspective of studying LLMs’ disadvantages.
+    We investigate LLMs’ vulnerability for some special inputs in terms of challenging their token segmentation, which provides a new perspective of studying LLMs’ disadvantages.
   2. 2.
@@
   3. 3.
-    Our experiment results obviously reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses, which can shed light on the subsequent work of improving LLMs through optimizing their tokenization.
+    Our experiment results obviously reveal the relationship between LLMs’ unsatisfactory tokenization and inaccurate responses, which can shed light on subsequent work of improving LLMs through optimizing tokenization.
 2 Related Work
@@
 Tokenization is a basic but crucial step in NLP.
-In order to get the models to better understand the text, great improvements have been made to the tokenization process.
-Currently, the mainstream tokenization approachs are based on sub-word units, including Byte Pair Encoding (BPE), WordPiece, and Unigram.
-BPE sennrich2016neural forms an extensive vocabulary starting from the base of character-level tokens, merging these token pairs and intercalating them into the vocabulary.
-The merging rule is to select the adjacent pairs with the highest word frequency.
-LLMs including GPT-3 brown2020language , RoBERTa liu2019roberta , BART lewis2019bart , Llama2 touvron2023llama , and Baichuan yang2023baichuan are based on BPE.
-WorldPiece Schuster_Nakajima_2012 is similar to BPE, but it differs in the strategy of merging pairs with reference to mutual information rather than frequency.
-LLMs built upon WordPiece include BERT devlin2019bert , DistilBERT sanh2020distilbert , and Electra clark2020electra .
-Different from these two algorithms, Unigram kudo2018subword starts with a large vocabulary and gradually trims it down to a smaller one.
-It measures the importance of subwords by computing the loss incurred by deleting different subwords, retaining those with high significance.
-LLMs utilizing Unigram include AlBERT lan2020albert and mBART liu2020multilingual .
-Besides, many tokenization tools integrating these algorithms have been developed, such as Google’s SentencePiece kudo2018sentencepiece and OpenAI’s tiktoken [^3] https://github.com/openai/tiktoken, which simplify the process of LLMs tokenization greatly.
-However, there has been no work concerning the shared risks inherent in these algorithms.
-So we firstly focus on this issue, discussing the underlying risks of LLMs’ tokenization.
+Currently, the mainstream tokenization approachs are based on subwords, including Byte Pair Encoding (BPE), WordPiece and Unigram.
+BPE Sennrich et al.
+(2016) forms the vocabulary starting from character-level tokens, merging token pairs and intercalating them into vocabulary.
+The merging rule is to select adjacent pairs with the highest word frequency.
+LLMs including GPT-3 Brown et al.
+(2020), RoBERTa Liu et al.
+(2019), and Llama2 Touvron et al.
+(2023) are based on BPE.
+WordPiece Schuster and Nakajima (2012) is similar to BPE, but it differs in the strategy of merging pairs with reference to mutual information rather than frequency.
+LLMs built upon WordPiece include BERT Devlin et al.
+(2019), DistilBERT Sanh et al.
+(2019), and Electra Clark et al.
+(2020).
+Different from these two algorithms, Unigram Kudo (2018) starts with a large vocabulary and gradually trims it down to a smaller one.
+It measures the importance of subwords by calculating loss associated with the removal of each subword, ultimately retaining those exhibit high importance.
+LLMs utilizing Unigram include AlBERT Lan et al.
+(2020) and mBART Liu et al.
+(2020).
+Besides, many tokenization tools integrating these algorithms have been developed, such as Google’s SentencePiece Kudo and Richardson (2018) and OpenAI’s tiktoken [^1] https://github.com/openai/tiktoken, which simplify the process of LLMs tokenization greatly.
+In recent years, it has been investigated how the input segmentation of pre-trained language models (PLMs) affects the interpretations of derivationally complex English words Hofmann et al.
+(2021).
+Some scholars have proposed FLOTA Hofmann et al.
+(2022), a simple yet effective method to improve PLMs’ tokenization of English words.
+However, there has been no work concerning the shared risks of LLMs in terms of inaccurate tokenization result.
+We focus on this issue, discussing the underlying risks of LLMs’ tokenization.
 Attack techniques in LLMs.
 With the growing prominence of LLMs, the security and vulnerability of these models have attracted significant attention, and even advanced LLMs like GPT-4 are no exception.
-A surge of research in this field is underway, with researchers launching attacks on LLMs from various aspects esmradi2023comprehensive ; chowdhury2024breaking including but not limited to, Prompt Injection choi2022prompt , Model Theft krishna2020thieves , Data Reconstruction carlini2021extracting , Data Poisoningwallace2021concealed ; xu2024instructions , and Member Inference Attack 7958568 .
-For instance, Prompt Injection Attack refers to a scenario where an attacker crafts malicious prompts with the intention of deceiving a language model to generate outputs inconsistent with its training data and anticipated functionality.
-Associating with PI attacks, threat actors aim for information gathering, fraud, intrusion, malware, content manipulation, and availability attacks choi2022prompt .
-Carlini, Nicholas, et al carlini2021extracting executed the Data Reconstruction attack on GPT-2, extracting personal identity information, code, and UUIDs.
+A surge of research in this field is underway, with researchers launching attacks on LLMs from various aspects Esmradi et al.
+(2023); Chowdhury et al.
+(2024) including but not limited to, Prompt Injection Choi et al.
+(2022), Model Theft Krishna et al.
+(2020), Data Reconstruction Carlini et al.
+(2021), Data PoisoningWallace et al.
+(2021); Xu et al.
+(2023), and Member Inference Attack Liu et al.
+(2023).
+For instance, Prompt Injection Attack refers to a scenario where an attacker crafts malicious prompts to deceive language models into generating outputs inconsistent with their training data and anticipated functionality.
+Threat actors aim at information gathering, fraud, intrusion, content manipulation, and availability attacks Choi et al.
+(2022).
+Carlini, Nicholas, et al Carlini et al.
+(2021) executed the Data Reconstruction attack on GPT-2, extracting personal identity information, code, and UUIDs.
 Data Poisoning pertains to the deliberate introduction of corrupted or malicious data into the training dataset to manipulate the model’s behavior.
-Member Inference Attack indicates that models tend to memorize sensitive information, posing them as potential targets for attacks.
-A study postulates that an attacker can train a binary classifier using the input-output pairs from LLMs to determine whether a specific individual was part of the model’s training set 7958568 .
-Diverging from existing works, our research innovatively suggests attacking the capabilities of language models from the perspective of Tokenization.
+Diverging from existing works, our research innovatively suggests attacking the capabilities of LLMs from the perspective of tokenization.
 3 Methodology of Dataset Construction
 In this section, we describe the process of constructing ADT (Adversarial Dataset for Tokenizer) in detail, which is used to challenge LLMs’ tokenization and thus reveal LLMs’ vulnerability.
-Our ADT contains two subsets, the manually constructed ADT-Human and automatically generated ADT-Auto.
+ADT contains two subsets, manually constructed ADT-Human and automatically generated ADT-Auto.
 3.1 Vocabulary Export
 In fact, an instance in our dataset comprises one sentence containing the token (word) that could challenge LLMs’ tokenization and one question related to the sentence.
-To inspect whether LLMs can accurately recognize these challenging tokens within various contexts based on their memories, the tokens should come from the vocabularies of LLMs.
-Therefore, the first step of our dataset construction is to export the vocabularies of LLMs.
-Given that Chinese is more complex and challenging than English in terms of tokenization, as detailed in Section 3.2, we pay more attention to the issues in Chinese data.
-Specifically, we select five widely-used open-source LLMs which are trained on extensive Chinese corpus, to export their Chinese vocabularies, including Chatglm3-6B zeng2023glm130b , Baichuan2-13B-Chat yang2023baichuan , Yi-34B-Chat ai2024yi , Qwen-7B-Chat bai2023qwen , and Qwen1.5-72B-Chat bai2023qwen .
-Besides, we export the English vocabularies from three English-based LLMs, i.e., Llama-3-8B-Instruct touvron2023llama , Llama-3-70B-Instruct touvron2023llama , and Mixtral-8x7B-Instruct-v0.1 jiang2024mixtral .
+To inspect whether LLMs can accurately recognize these challenging tokens within various contexts based on their memories, the tokens should come from vocabularies of LLMs.
+Therefore, the first step of dataset construction is to export vocabularies of LLMs.
+Given that Chinese is more complex and challenging than English in terms of tokenization, as detailed in Section 3.2, we pay more attention to the issue in Chinese data.
+Specifically, we select five widely-used open-source LLMs which are trained on extensive Chinese corpus, to export their Chinese vocabularies, including Chatglm3-6B Zeng et al.
+(2023), Baichuan2-13B-Chat Yang et al.
+(2023), Yi-34B-Chat Young et al.
+(2024), Qwen-7B-Chat Bai et al.
+(2023), and Qwen1.5-72B-Chat Bai et al.
+(2023).
+Besides, we export the English vocabularies from three English-based LLMs, i.e., Llama-3-8B-Instruct Touvron et al.
+(2023), Llama-3-70B-Instruct Touvron et al.
+(2023), and Mixtral-8x7B-Instruct-v0.1 Jiang et al.
+(2024).
 Notably, the Chinese vocabulary of Qwen-7B-Chat and Qwen1.5-72B-Chat is the same, so is the English vocabulary of Llama-3-8B-Instruct and Llama-3-70B-Instruct.
-These LLMs use SentencePiece or tiktoken to train their tokenizers, which are tokenizer tools widely used by various language models.
-In the process of exporting the vocabularies, a sequence of straightforward operations is considered.
-Initially, the tokenizer is decoded to obtain the vocabulary, followed by the removal of leading and trailing spaces from each token.
-Notably, if SentencePiece is used for tokenization in the training phase, some certain tokens may begin with a special token ‘_’ because SentencePiece treats the input text just as a sequence of Unicode characters, and whitespace is also handled as a normal symbol.
-To handle the whitespace as a basic token explicitly, SentencePiece first escapes the whitespace with a meta symbol ‘_’ (U+2581)  [^4] https://github.com/google/sentencepiece.
-Consequently, when exporting the vocabularies for models trained with SentencePiece, we replace ‘_’ with a blank character.
+Language Model Vocabulary Size Vocabulary Size of Specific Language Chinese Chatglm3-6B 64,789 30,922 Baichuan2-13B-Chat 125,696 70,394 Yi-34B-Chat 64,000 21,353 Qwen-7B-Chat 151,851 24,953 Qwen1.5-72B-Chat 151,646 24,953 English Llama-3-8B-Instruct 128,257 72,420 Llama-3-70B-Instruct 128,257 72,420 Mixtral-8x7B-Instruct-v0.1 32,000 25,056
+Table 1: Vocabulary sizes of different LLMs.
+In the process of exporting vocabularies, a sequence of straightforward operations is considered.
+Initially, the tokenizer is decoded to obtain vocabulary, followed by the removal of leading and trailing spaces from each token.
+Notably, if SentencePiece is used for tokenization in training phase, some certain tokens may begin with a special token ‘_’ because SentencePiece treats the input text just as a sequence of Unicode characters, and whitespace is also handled as a normal symbol.
+To handle the whitespace as a basic token explicitly, SentencePiece first escapes the whitespace with a meta symbol ‘_’ (U+2581) Kudo and Richardson (2018).
+Consequently, when exporting vocabularies for models trained with SentencePiece, we replace ‘_’ with a blank character.
 We summarize the exported vocabulary sizes of different LLMs in Table 1.
-Table 1: Vocabulary sizes of different LLMs.
-Language Model Vocabulary Size Vocabulary Size
-of Specific Language
-Chinese Chatglm3-6B 64,789 30,922
-Baichuan2-13B-Chat 125,696 70,394
-Yi-34B-Chat 64,000 21,353
-Qwen-7B-Chat 151,851 24,953
-Qwen1.5-72B-Chat 151,646 24,953
-English Llama-3-8B-Instruct 128,257 72,420
-Llama-3-70B-Instruct 128,257 72,420
-Mixtral-8x7B-Instruct-v0.1 32,000 25,056
 3.2 ADT-Human Construction
-Based on the exported vocabularies, we manually construct our dataset ADT-Human, to challenge and evaluate the tokenization of different LLMs.
-The process of manual construction is depicted in Figure 2.
+Based on exported vocabularies, we manually construct dataset ADT-Human, to challenge and evaluate the tokenization of different LLMs. The process of manual construction is depicted in Figure 2.
 Figure 2: Our framework of constructing ADT-Human manually.
-Specifically, we sophisticatedly select eleven tokens from each Chinese vocabulary, and seven tokens for each English vocabulary.
+Approach Origin token After insertion Model tokenization Human tokenization (Challenging span) Before Schema BC ABC [A, BC] [AB, C] Example stable movestable [move, stable] [moves, table] After Schema AB ABC [AB, C] [A, BC] Example driven drivenothing [driven, othing] [drive, nothing] Before & After Schema BC ABCD [A, BC, D] [AB, CD] Example this sixthisland [six, this, land] [sixth, island]
+Table 2: Three approaches of generating challenging spans.
+Our purpose is to confirm the existence of challenges in tokenization, so for each LLM, a certain amount of data is constructed, which does not need to be large.
+Specifically, we select eleven tokens from each Chinese vocabulary, and seven tokens for each English vocabulary.
+The main criterion for selecting tokens is that they should be easy to make sentences with.
+According to the experimental results presented in Section 4.2, these data can effectively challenge the performance of LLMs, proving this issue deserves more attention.
+As for efficiently generating data in bulk, we design the framework for automatic generation in Section 3.3.
 Then, for each selected token, we adopt one of the three approaches listed in Table 2, to convert it into a challenging span through inserting a special character span before or (and) after it.
 These challenging spans would disrupt the conventional tokenization process, thus confusing LLMs.
+The schemas and examples of three approaches are also shown in the second and third parts of Figure 2.
 The considerations in the three conversion approaches are introduced as follows.
   1. 1.
-    Before: A character span $s$ is inserted before the origin token, causing that the concatenation of $s$ and the token’s prefix is just another token (word) existing in the vocabulary, as ‘move’+‘s’$\rightarrow$‘moves’ in Table 2.
+    Before: A character span $s$ is inserted before origin token, causing the concatenation of $s$ and the token’s prefix is just another token existing in vocabulary, as ‘move’+‘s’$\rightarrow$‘moves’ in Table 2.
   2. 2.
-    After: A character span $s$ is inserted after the origin token, causing that the concatenation of the token’s suffix and $s$ is just another token existing in the vocabulary, as ‘n’+‘othing’$\rightarrow$‘nothing’ in Table 2.
+    After: A character span $s$ is inserted after origin token, causing the concatenation of the token’s suffix and $s$ is just another token existing in vocabulary, as ‘n’+‘othing’$\rightarrow$‘nothing’ in Table 2.
   3. 3.
-    Before & After: Character spans $s_{1}$ and $s_{2}$ are inserted before and after the origin token, respectively, causing that the concatenation of $s_{1}$ and the token’s prefix is just another token existing in the vocabulary, meanwhile so is the concatenation of the token’s suffix and $s_{2}$, as ‘six’+‘th’$\rightarrow$‘sixth’ and ‘is’+‘land’$\rightarrow$‘island’ in Table 2.
+    Before & After: Character spans $s_{1}$ and $s_{2}$ are inserted before and after origin token, respectively, causing the concatenation of $s_{1}$ and the token’s prefix is just another token existing in vocabulary, meanwhile so is the concatenation of the token’s suffix and $s_{2}$, as ‘six’+‘th’$\rightarrow$‘sixth’ and ‘is’+‘land’$\rightarrow$‘island’ in Table 2.
 Next, for each challenging span $s_{c}$, we manually compose a corresponding instance as the data in ADT-Human.
 One instance consists of a sentence in which $s_{c}$ presents, and a corresponding question of which the answer comes from $s_{c}$.
-In Appendix A, we list all instances (along with their correct tokenization) in ADT-Human for the four Chinese vocabularies and two English vocabularies we export.
-As we can see, it is challenging for the LLMs to understand the instances due to the presence of challenging spans.
-Table 2: Three approaches of generating challenging spans.
-Approach Origin token After insertion Model tokenization Human tokenization
-(Challenging span)
-Before Schema BC ABC [A, BC] [AB, C]
-Example stable movestable [move, stable] [moves, table]
-After Schema AB ABC [AB, C] [A, BC]
-Example driven drivenothing [driven, othing] [drive, nothing]
-Before & After Schema BC ABCD [A, BC, D] [AB, CD]
-Example this sixthisland [six, this, land] [sixth, island]
+In Appendix A, we list all instances (along with their correct tokenizations) of ADT-Human.
+As we can see, it is challenging for LLMs to understand the instances due to the presence of challenging spans.
 It’s worth noting that the tokenization difficulty of English is less than that of Chinese, since spaces are regularly used as delimiters to separate each word from others in English.
@@
 Thus, during the manual construction of English instances, we deliberately exclude the spaces between some tokens to provoke challenges to tokenization process.
+This decision stems from the recognition that powerful models should possess robust abilities across various scenarios which can occur in real-world applications, including handling cases where spaces might be omitted in English text input Hofmann et al.
+(2022).
 3.3 ADT-Auto Generation
-Given the inefficiency of constructing dataset manually, we further develop an automatic construction framework for the dataset to challenge LLMs’ tokenization.
+Figure 3: Our framework of generating ADT-Auto automatically.
+Given the inefficiency of constructing dataset manually, we further develop an automatic generation framework for dataset to challenge LLMs’ tokenization.
 As discussed in Section 3.2, the challenges of tokenization in English are less severe than in Chinese.
-Consequently, we primarily concentrate on the automatic generation of dataset in Chinese.
+Consequently, we primarily concentrate on the automatic generation of Chinese data.
 Figure 3 illustrates the process of automatically constructing our dataset ADT-Auto.
-Figure 3: Our framework of generating ADT-Auto automatically.
 3.3.1 Word-pair Matching
-The automatic generation of dataset is also based on the exported vocabularies.
+The automatic generation of dataset is also based on exported vocabularies.
 From the vocabularies, we first seek some qualified word pairs.
-Given two words Word 1 and Word 2, they are considered to be match when meeting the following criteria: The suffix (or whole) of Word 1 can be concatenated with the prefix (or whole) of Word 2, as a token (word) existing in the vocabulary, denoted as Trap Word.
+Given two words Word 1 and Word 2, they are considered to be match when meeting the following criteria: The suffix (or whole) of Word 1 can be concatenated with the prefix (or whole) of Word 2, as a token existing in the vocabulary, denoted as Trap Word.
 Meanwhile, the concatenation of Word 1 and Word 2 should not be a token existing in the vocabulary.
@@
 Above clarifications indicate that the concatenation of Word 1 and Word 2 (as ‘ABCD’ in Figure 3) corresponds to the challenging span in the construction process of ADT-Human.
-Our objective is to compose a sentence that can not only convey the semantic essence of both Word 1 and Word 2, but also induce LLM to isolate Trap Word during the tokenization process.
-Notably, we neglect the situation that the concatenation of Word 1 and Word 2 is just Trap Word, because it would not pose challenges to LLMs’ tokenization.
-To augment the efficacy of Trap Word upon matching, we also consider the word-pair matching criterion that the remaining parts of Word 1 and Word 2 excluded by Trap Word should also exist in the vocabulary as a token (word).
+Our goal is to compose a sentence that not only convey the semantic essence of Word 1 and Word 2, but also induce LLM to isolate Trap Word when tokenization.
+Notably, we ignore the situation that the concatenation of Word 1 and Word 2 is just Trap Word, as it would not pose challenges to LLMs’ tokenization.
+To augment the matching efficacy of Trap Word upon matching, we also consider the criterion that the remaining parts of Word 1 and Word 2 excluded by Trap Word should also exist in the vocabulary as a token.
 Furthermore, the first or last character of Trap Word cannot be a Chinese stop character.
 3.3.2 Instance Generalization
-With each Word 1, Word 2 and their corresponding Trap Word, we harness GPT-4 to generate an instance of ADT-Auto, which consists of a sentence and a corresponding question.
+With each Word 1, Word 2 and their corresponding Trap Word, we harness GPT-4 to generate an instance of ADT-Auto, which consists of a sentence and a question.
 Specifically, we prompt GPT-4 to generate a sentence that is used to challenge the tokenization of LLMs.
@@
 In addition, GPT-4 is also required to devise a question related to the generated sentence, which is used to evaluate LLMs’ tokenization performance through their answers.
-The instance must meet the criteria as below: The sentence should convey the meanings of both Word 1 and Word 2, while the question’s answer should come from Word 1, Word 2, Trap Word or their combination.
-Thus, the influence of LLMs’ incorrect tokenization can be identified by checking their answers to the question.
-The prompt for GPT-4 to generate ADT-Auto instances also includes some demonstration examples in addition to the task instructions, which is illustrated in Appendix B.
+The instance must meet criteria as below: The sentence should convey the meanings of both Word 1 and Word 2, while the question’s answer should come from Word 1, Word 2, Trap Word or their combination.
+Thus, the influence of LLMs’ incorrect tokenization can be identified by checking answers to the question.
+The prompt for GPT-4 to generate instances includes some demonstration examples in addition to task instruction, which is illustrated in Appendix D.
 3.3.3 Filtering Based on Trap Word
 The goal of our dataset is to expose the flaw of LLMs’ tokenization.
-Therefore, we will only retain the instances that can induce the tokenization problem of LLMs.
-To this end, we check each instance generated at the previous step.
+Therefore, we will only retain the instances that can induce tokenization problem of LLMs, so we check each instance generated at the previous step.
 For these instances, the presence of Trap Word implies a challenging case that is likely to induce tokenization problems.
-Accordingly, we first select an open-source LLM and get the Trap Word set based on its vocabulary.
-Given an instance, we retain it if a certain Trap Word is both in the LLM’s tokenization list for the instance’s sentence (as List 1 in Figure 3) and in its tokenization list for the answer (as List 2 in Figure 3).
+Given an instance, we retain it if its corresponding Trap Word is both in the LLM’s tokenization list for the instance’s sentence (as List 1 in Figure 3) and in its tokenization list for the answer (as List 2 in Figure 3).
 Such filtering criterion indicates that the LLM commits tokenization errors on understanding the sentence and response for the instance.
-Thus, this instance is also likely to challenge other LLMs’ tokenization.
 3.3.4 Manual Filtering
-To ensure that the retained instances can induce tokenization problems and meanwhile have reasonable expressions, we further adopt the manual assessment for instances.
-Specifically, we further select the high-qualify instances with considering their sentence expressions and LLMs’ responses.
+To ensure the retained instances can induce tokenization problems and meanwhile have reasonable expressions, we further adopt manual assessment for instances.
+Specifically, we select the high-qualify instances with considering sentence expressions and LLM’s responses.
 Notably, we might still retain some instances to which the used LLM has correct response, since the other LLMs are still likely to commit inaccurate tokenization for these instances, resulting in unsatisfactory responses.
-Due to space limitation, we take Qwen-7B as an example to illustrate the process of generating ADT-Auto instances.
+Due to space limitation, we take Qwen-7B as an example to illustrate the process of generating instances.
 There are 24,953 Chinese tokens in Qwen-7B’s vocabulary, and after word-pair matching, 1,764,692 word-pairs are obtained.
-From the matching word-pairs, 8,000 pairs are selected randomly and then used for instance generation by GPT-4.
+From the matching word-pairs, 8,000 pairs are selected randomly and used for instance generation by GPT-4.
 Due to the inherent stochastic characteristic of LLMs on response generation, we conduct three iterations of filtering operations introduced in Subsection 3.3.3 to get more qualified instances, and thus retain 894 instances.
-Next, through the manual filtering introduced in Subsection 3.3.4, we retain 231 for ADT-Auto in the end.
+Next, through the manual filtering introduced in Subsection 3.3.4, we retain 231 instances for ADT-Auto in the end.
 4 Experiments
@@
 Considering the open-source LLMs used in our experiments, we select the LLMs previously used in the construction of ADT-Human, including Chatglm3-6B, Baichuan2-13B-Chat, Yi-34B-Chat, Qwen-7B-Chat, and Qwen1.5-72B-Chat for Chinese data, as well as Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1 for English data.
-We test these LLMs using both locally deployed versions and API versions.
-For the closed-source LLMs used in our experiments, we test GPT-4o, GPT-4, GPT-3.5-Turbo, Qwen2.5-max, step-1-8k [^5] https://platform.stepfun.com/docs/llm/text, moonshot-v1-8k [^6] https://platform.moonshot.cn/docs, ERNIE-3.5-8K777 https://cloud.baidu.com/doc/WENXINWORKSHOP/s/jlil56u11 for Chinese data, and GPT-4o, GPT-4 and GPT-3.5-Turbo for English data.
+We test these LLMs using both locally deployed versions and API versions, with the exception of Chatglm3-6B, which does not have API version.
+For the closed-source LLMs, we test GPT-4o, GPT-4, GPT-3.5-Turbo, Qwen2.5-max, step-1-8k [^2] https://platform.stepfun.com/docs/llm/text, moonshot-v1-8k [^3] https://platform.moonshot.cn/docs, ERNIE-3.5-8K444 https://cloud.baidu.com/doc/WENXINWORKSHOP/s/jlil56u11 for Chinese data, and GPT-4o, GPT-4 and GPT-3.5-Turbo for English data.
+Additionally, we test the API version of Deepseek-R1, which has recently gained significant attention, for Chinese data.
 In our experiments, we directly use the dataset ADT constructed with the method introduced in Section 3, which includes the manually constructed ADT-Human (containing Chinese and English instances) and the automatically generated ADT-Auto (only containing Chinese instances).
@@
 4.2 ADT-Human’s Challenges to LLMs
-Firstly, we investigate the challenges posed by our manually constructed dataset ADT-Human to LLMs.
-To this end, we evaluate the experimented LLMs’ performance on the dataset, through counting the number of incorrect answers generated by the LLMs for the questions in the instances.
-Recalling the process of manually composing challenging spans introduced in Section 3.2, the span ‘BC’ in Table 2 is in fact a Trap Word mentioned in Section 3.3.
-Hence, for a given LLM, its answer including a Trap Word is identified as its inaccurate response for the question undoubtedly.
-Such a case also indicates that the LLM generates a wrong token segmentation.
-The percentage of incorrect responses provided by the LLMs for the Chinese data (corresponding to the four LLMs’ vocabularies) and English data (corresponding to the two LLMs’ vocabularies) are presented in Table 3 and Table 4, respectively.
-It should be noted that there is no API version of Chatglm3-6B.
-The results shows that the ADT-Human poses a significant challenge to both open-source and closed-source LLMs on their tokenization, resulting in very high rates of inaccurate responses.
+Firstly, we investigate the challenges posed by the manually constructed dataset ADT-Human to LLMs.
+Specifically, we evaluate LLMs’ performance through counting the number of incorrect answers generated by LLMs for the questions in instances.
+Recalling the process of manually composing challenging spans introduced in Section 3.2, the span ‘BC’ in Table 2 is in fact the Trap Word mentioned in Section 3.3.
+Hence, for a given LLM, its answer including a Trap Word is identified as inaccurate response for the question undoubtedly.
+We identify the correctness of LLMs’ responses through human assessment.
+The percentage of incorrect responses provided by the LLMs for Chinese data (corresponding to the four LLMs’ vocabularies) and English data (corresponding to the two LLMs’ vocabularies) are presented in Table 3 and Table 4, respectively.
+The results show that ADT-Human poses a significant challenge to both open-source and closed-source LLMs on their tokenization, resulting in very high rates of inaccurate responses.
+It is also worth mentioning that the recent state-of-the-art (SOTA) GPT-4o cannot outperform GPT-4 , implying that the advancements of these LLMs have not yet addressed this primary but challenging problem.
+Model Source LLM of vocabulary Overall error rate Chatglm3 Baichuan2 Yi Qwen Open-source (Local) Chatglm3-6B 100.00 100.00 100.00 90.91 97.73 Baichuan2-13B-Chat 90.91 100.00 81.82 100.00 93.18 Yi-34B-Chat 72.73 63.64 100.00 100.00 84.09 Qwen-7B-Chat 100.00 72.73 90.91 100.00 90.91 Qwen1.5-72B-Chat 90.91 45.45 81.82 100.00 79.55 Open-source (API) Baichuan2-13B-Chat 100.00 100.00 90.91 100.00 97.73 Yi-34B-Chat 81.82 54.55 100.00 90.91 81.82 Qwen-7B-Chat 100.00 81.82 81.82 100.00 90.91 Qwen1.5-72B-Chat 90.91 54.55 81.82 100.00 81.82 DeepSeek-R1 36.36 27.27 45.45 54.55 40.91 Closed-source GPT-4o 72.73 27.27 54.55 90.91 61.36 GPT-4 81.82 45.45 45.45 27.28 50.00 GPT-3.5-Turbo 72.73 27.27 72.73 72.73 61.36 Qwen2.5-max 90.91 72.73 90.91 100.00 88.64 step-1-8k 63.64 18.18 72.73 63.64 54.55 moonshot-v1-8k 81.82 27.27 100.00 81.82 72.73 ERNIE-3.5-8K 72.73 54.55 54.55 72.73 63.64
 Table 3: Error rates (%) of answers on ADT-Human (Chinese).
-Model Source LLM of vocabulary Overall error rate Chatglm3 Baichuan2 Yi Qwen Open-source (Local) Chatglm3-6B 100.00 100.00 100.00 90.91 97.73 Baichuan2-13B-Chat 90.91 100.00 81.82 100.00 93.18 Yi-34B-Chat 72.73 63.64 100.00 100.00 84.09 Qwen-7B-Chat 100.00 72.73 90.91 100.00 90.91 Qwen1.5-72B-Chat 90.91 45.45 81.82 100.00 79.55 Open-source (API) Baichuan2-13B-Chat 100.00 100.00 90.91 100.00 97.73 Yi-34B-Chat 81.82 54.55 100.00 90.91 81.82 Qwen-7B-Chat 100.00 81.82 81.82 100.00 90.91 Qwen1.5-72B-Chat 90.91 54.55 81.82 100.00 81.82 Closed-source GPT-4o 72.73 27.27 54.55 90.91 61.36 GPT-4 81.82 45.45 45.45 27.28 50.00 GPT-3.5-Turbo 72.73 27.27 72.73 72.73 61.36 Qwen2.5-max 90.91 72.73 90.91 100.00 88.64 step-1-8k 63.64 18.18 72.73 63.64 54.55 moonshot-v1-8k 81.82 27.27 100.00 81.82 72.73 ERNIE-3.5-8K 72.73 54.55 54.55 72.73 63.64
+Model Source LLM of vocabulary Overall error rate Llama-3 Mixtral Open-source (Local) Llama-3-8B-Instruct 100.00 85.71 92.86 Llama-3-70B-Instruct 57.14 71.43 64.29 Mixtral-8x7B-Instruct-v0.1 85.71 100.00 92.86 Open-source (API) Llama-3-8B-Instruct 100.00 71.43 85.71 Llama-3-70B-Instruct 57.14 28.57 42.86 Mixtral-8x7B-Instruct-v0.1 71.43 100.00 85.71 Closed-source GPT-4o 57.14 71.43 64.29 GPT-4 57.14 57.14 57.14 GPT-3.5-Turbo 71.43 57.14 64.29
 Table 4: Error rates (%) of answers on ADT-Human (English).
-Model Source LLM of vocabulary Overall error rate Llama-3 Mixtral Open-source (Local) Llama-3-8B-Instruct 100.00 85.71 92.86 Llama-3-70B-Instruct 57.14 71.43 64.29 Mixtral-8x7B-Instruct-v0.1 85.71 100.00 92.86 Open-source (API) Llama-3-8B-Instruct 100.00 71.43 85.71 Llama-3-70B-Instruct 57.14 28.57 42.86 Mixtral-8x7B-Instruct-v0.1 71.43 100.00 85.71 Closed-source GPT-4o 57.14 71.43 64.29 GPT-4 57.14 57.14 57.14 GPT-3.5-Turbo 71.43 57.14 64.29
-It is also worth mentioning that the recent state-of-the-art (SOTA) GPT-4o cannot outperform GPT-4 (higher error rate), implying that the advancements of these LLMs have not yet addressed this primary but challenging problem.
+To further investigate the impact of tokenization on the performance of LLMs, a fine-grained analysis of the relationship between tokenization and response is conducted.
+In Appendix B, we quantitatively examine the relationship between the correctness of tokenization and the correctness of LLM’s response.
+Given the inaccessibility of tokenization lists for closed-source LLMs, we perform a statistical analysis on open-source LLMs.
+Since LLMs’ API versions do not directly provide tokenization lists, the tokenization results obtained from the corresponding locally deployed versions are used.
+Appendix B.1 and Appendix B.2 respectively illustrate the quantitative relationships between tokenization and response for the LLMs tested on Chinese and English data of ADT-Human.
+Right Wrong
+response response
+Right tokenization 3 3
+Wrong tokenization 1 37
+Figure 4: Four relationships between tokenization and response, take Qwen-7B-Chat as an example.
+This study further intuitively illustrates the relationships between tokenization and response in the form of pie charts in Appendix C.
+In Appendix C.1 and Appendix C.2, the proportions of four relationships between tokenization and response for each LLM on Chinese and English data of ADT-Human are displayed respectively.
+Figure 4 illustrates the quantitative relationships and pie chart using Qwen-7B-Chat as an example.
+For more details, please refer to Appendix B and Appendix C.
+We mainly focus on the proportion of TN (tokenization incorrect and response incorrect).
+As shown in pie charts, the proportion of TN cases is very high in ADT-Human, with an average of 80.91% for Chinese data and 79.78% for English data.
+This indicates that tokenization errors significantly affect the accuracy of LLM responses and also demonstrates that ADT-Human effectively challenges the performance of LLMs.
 4.3 ADT-Auto’s Challenges to LLMs
 Similar to the investigation of ADT-Human’s challenges to LLMs’ tokenization, we also evaluate the LLMs’ performance on ADT-Auto.
-We still test five open-source LLMs with their both locally deployed and API versions, and seven closed-source LLMs only having APIs.
+Model Fraction Error rate (%) Open-source (Local) Chatglm3-6B 156/231 67.53 Baichuan2-13B-Chat 147/231 63.64 Yi-34B-Chat 90/231 38.96 Qwen-7B-Chat 185/231 80.09 Qwen1.5-72B-Chat 93/231 40.26 Open-source (API) Baichuan2-13B-Chat 167/231 72.29 Yi-34B-Chat 80/231 34.63 Qwen-7B-Chat 160/231 69.26 Qwen1.5-72B-Chat 97/231 41.99 DeepSeek-R1 57/231 24.68 Closed-source GPT-4o 75/231 32.47 GPT-4 89/231 38.53 GPT-3.5-Turbo 99/231 42.86 Qwen2.5-max 75/231 32.47 step-1-8k 60/231 25.97 moonshot-v1-8k 60/231 25.97 ERNIE-3.5-8K 51/231 22.08
+Table 5: Error rates of answers on ADT-Auto.
 ADT-Auto’s instances come from Qwen-7B’s vocabulary, and the rates of LLMs’ inaccurate responses for these instances’ questions are listed in Table 5.
 Based on the results, we have the following observations and analysis:
-Table 5: Error rates of answers on ADT-Auto.
-Model Fraction Error rate(%)
-Open-source (Local) Chatglm3-6B 156/231 67.53
-Baichuan2-13B-Chat 147/231 63.64
-Yi-34B-Chat 90/231 38.96
-Qwen-7B-Chat 185/231 80.09
-Qwen1.5-72B-Chat 93/231 40.26
-Open-source (API) Baichuan2-13B-Chat 167/231 72.29
-Yi-34B-Chat 80/231 34.63
-Qwen-7B-Chat 160/231 69.26
-Qwen1.5-72B-Chat 97/231 41.99
-Closed-source GPT-4o 75/231 32.47
-GPT-4 89/231 38.53
-GPT-3.5-Turbo 99/231 42.86
-Qwen2.5-max 75/231 32.47
-step-1-8k 60/231 25.97
-moonshot-v1-8k 60/231 25.97
-ERNIE-3.5-8K 51/231 22.08
   1. 1.
-    Compared with the closed-source LLMs, the open-source LLMs suffer from ADT-Auto’s challenges more apparently.
-It implies that these closed-source LLMs, as the profit-making flagships of their creator companies, naturally have stronger capabilities than the open-source LLMs that are created for public usage.
+    Compared with closed-source LLMs, open-source LLMs suffer from ADT-Auto’s challenges more apparently.
+It implies that these closed-source LLMs, as the profit-making flagships of their creator companies, naturally have stronger capabilities than open-source LLMs that are created for public usage.
   2. 2.
@@
   3. 3.
-    Qwen1.5-72B-Chat has much lower error rates than Qwen-7B-Chat, although they have the same vocabulary.
-We specially check the instances to which Qwen1.5-72B-Chat gives the correct answers but Qwen-7B-Chat gives the wrong answers, and find the two models have the same incorrect tokenization lists for these instances.
-It suggests that their different performance is not caused by their tokenization.
-These results also imply that in the case of incorrect tokenization, the bigger LLMs are more robust and likely to generate correct responses than the smaller LLMs thanks to their stronger capabilities brought by the larger scale.
-5 Perspectives and conclusion
-Based on our findings in this paper, for the work dedicating to improve LLMs’ tokenization, and so as to improve LLMs’ performance on varied tasks, it is reasonable and promising to expand LLMs’ vocabulary sizes or developing more powerful tokenizers or effective tokenization algorithms.
-As the idea proposed by Kudo kudo2018subword , employing multiple subword segmentation methods might be a feasible strategy to enhance the robustness of LLMs’ tokenization process.
-We hope that our work in this paper and the dataset could shed light on the subsequent research on improving LLMs’ tokenization and then the performance on other tasks.
+    Qwen1.5-72B-Chat has lower error rates than Qwen-7B-Chat, although they have the same vocabulary.
+We specially check some instances to which Qwen1.5-72B-Chat gives correct answers but Qwen-7B-Chat gives wrong answers, and find the two models have the same incorrect tokenization lists for these instances.
+It suggests that their different performance is not caused by tokenization.
+The results also imply that in the case of incorrect tokenization, the bigger LLMs are more robust and likely to generate correct responses than the smaller LLMs thanks to their stronger capabilities brought by the larger scale.
+Similar to Section 4.2, Appendix B.3 presents the quantitative relationships between tokenization and response for each open-source LLM tested on ADT-Auto.
+The corresponding pie charts are shown in Appendix C.3.
+As indicated by the pie charts, the proportion of TN cases is also very high on ADT-Auto, with an average of 46.11%.
+This further demonstrates that tokenization errors significantly impact the accuracy of LLMs’ responses and highlights the effectiveness of ADT-Auto in challenging the performance of LLMs.
+5 Conclusion
 In this paper, we dedicate to deeply investigating the relationship between LLMs’ vulnerability on tokenization and their unsatisfactory responses for some tasks.
-To this end, we construct an adversarial dataset ADT (Adversarial Dataset for Tokenizer) consisting of a manual constructed subset ADT-Human and an automatically generated subset ADT-Auto, of which each instance includes a sentence and a corresponding question.
-Our experiments obviously demonstrate that our dataset does challenge the studied open-source and closed-source LLMs’ token segmentation, resulting in their incorrect answers to the instances’ questions.
+To this end, we construct an adversarial dataset ADT (Adversarial Dataset for Tokenizer) consisting of a manually constructed subset ADT-Human and an automatically generated subset ADT-Auto, of which each instance includes a sentence and a corresponding question.
+Our experiments demonstrate that our dataset does challenge the studied open-source and closed-source LLMs’ token segmentation, resulting in their incorrect answers.
+We hope our work and dataset could shed light on the subsequent research on improving LLMs’ performance.
+Limitations
+The key contribution of this study lies in drawing attention to the impact of tokenization on LLM’s performance and providing two frameworks for data generation.
+As for how to propose improvement strategies based on this phenomenon, we are currently in the process of exploration.
+Additionally, this study focuses primarily on Chinese.
+Whether other languages are similarly affected by tokenization remains to be further investigated.
+Ethics Statement
+We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
+Use of Human Annotations.
+Human annotations are only used in methodological research at the beginning of the work, to assist in analyzing the feasibility of the proposed solution.
+Annotators consented to the use of data for research purposes.
+We ensure the privacy of all annotators is protected throughout the annotation process, and all of them are adequately paid according to local standards.
+Risks.
+Synthetic data generated by LLMs may involve potential ethical risks regarding fairness and bias Bommasani et al.
+(2021); Blodgett et al.
+(2020), which results in further consideration when they are employed in downstream tasks.
+We asked our members for proofreading to refine the offensive and harmful data generated by GPT-4.
+Despite these considerations, there may still be some unsatisfactory data that goes unnoticed in our final dataset.
 Appendix A Details of ADT-Human
@@
 A.5 Llama-3
+  1. 1.
+    In the postal history course, today’s homework is to list|envelope|classes in history. What should I do?
+  2. 2.
+    As a Marathon lover, in the past he only ran|domestically, but now he also goes abroad. What behavior of him is discussed in the sentence?
+  3. 3.
+    As fans of Kenshi Yonezu, they sing|lemon in the auditorium. What song are they singing in the sentence?
+  4. 4.
+    I wanted to go to the beach, but|tons|of work piled up. How much work did I have?
+  5. 5.
+    We are measuring how fat|her|cat is to make sure her cat is healthy. What metric are we measuring?
+  6. 6.
+    This automotive shop mainly sells car|bonnets. What does the shop sell?
+  7. 7.
+    The researcher was disappointed to miss|ionization in the sample, which was crucial for the experiment’s success. What made the researcher disappointed?
 A.6 Mixtral
-Appendix B Details of prompt
+  1. 1.
+    The analyst emphasized the importance of tracking live|return to gauge real-time performance. What importance was emphasized by the analyst?
+  2. 2.
+    Many countries import|sports from another country. What do those countries import?
+  3. 3.
+    He moves|table to another place. What is he doing?
+  4. 4.
+    Those pants and leather shoes fitted|speakers very well. Who do those pants and leather shoes fit well?
+  5. 5.
+    The soccer team won|derby against its rival. Did the soccer team win or lose?
+  6. 6.
+    They swap|pears with each other. What are they exchanging?
+  7. 7.
+    The leg|ends|up being the most injured part, requiring immediate medical attention. In the sentence, which part is injured the most?
+Appendix B The relationship between tokenization and response
+B.1 ADT-Human (Chinese)
+Right Wrong
+response response
+Right tokenization 2 1
+Wrong tokenization 1 40
+Table 6: (a) Baichuan2-13B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 1 2
+Wrong tokenization 0 41
+Table 7: (b) Baichuan2-13B-Chat (API)
+Right Wrong
+response response
+Right tokenization 3 0
+Wrong tokenization 5 36
+Table 8: (c) Yi-34B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 3 0
+Wrong tokenization 7 34
+Table 9: (d) Yi-34B-Chat (API)
+Right Wrong
+response response
+Right tokenization 3 3
+Wrong tokenization 1 37
+Table 10: (e) Qwen-7B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 3 3
+Wrong tokenization 1 37
+Table 11: (f) Qwen-7B-Chat (API)
+Right Wrong
+response response
+Right tokenization 6 0
+Wrong tokenization 3 35
+Table 12: (g) Qwen1.5-72B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 6 0
+Wrong tokenization 2 36
+Table 13: (h) Qwen1.5-72B-Chat (API)
+Right Wrong
+response response
+Right tokenization 1 1
+Wrong tokenization 0 42
+Table 14: (i) Chatglm3-6B (Local)
+Right Wrong
+response response
+Right tokenization 3 0
+Wrong tokenization 23 18
+Table 15: (j) Deepseek-R1 (API)
+B.2 ADT-Human (English)
+Right Wrong
+response response
+Right tokenization 1 0
+Wrong tokenization 0 13
+Table 16: (a) Llama-3-8B-Instruct (Local)
+Right Wrong
+response response
+Right tokenization 0 1
+Wrong tokenization 0 13
+Table 17: (b) Llama-3-8B-Instruct (API)
+Right Wrong
+response response
+Right tokenization 0 0
+Wrong tokenization 3 11
+Table 18: (c) Llama-3-70B-Instruct (Local)
+Right Wrong
+response response
+Right tokenization 0 0
+Wrong tokenization 8 6
+Table 19: (d) Llama-3-70B-Instruct (API)
+Right Wrong
+response response
+Right tokenization 0 0
+Wrong tokenization 2 12
+Table 20: (e) Mixtral-8x7B-Instruct-v0.1 (Local)
+Right Wrong
+response response
+Right tokenization 0 0
+Wrong tokenization 2 12
+Table 21: (f) Mixtral-8x7B-Instruct-v0.1 (API)
+B.3 ADT-Auto
+Right Wrong
+response response
+Right tokenization 28 32
+Wrong tokenization 56 115
+Table 22: (a) Baichuan2-13B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 23 37
+Wrong tokenization 41 130
+Table 23: (b) Baichuan2-13B-Chat (API)
+Right Wrong
+response response
+Right tokenization 45 14
+Wrong tokenization 96 76
+Table 24: (c) Yi-34B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 48 11
+Wrong tokenization 103 69
+Table 25: (d) Yi-34B-Chat (API)
+Right Wrong
+response response
+Right tokenization 4 22
+Wrong tokenization 42 163
+Table 26: (e) Qwen-7B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 10 16
+Wrong tokenization 61 144
+Table 27: (f) Qwen-7B-Chat (API)
+Right Wrong
+response response
+Right tokenization 21 5
+Wrong tokenization 117 88
+Table 28: (g) Qwen1.5-72B-Chat (Local)
+Right Wrong
+response response
+Right tokenization 20 6
+Wrong tokenization 114 91
+Table 29: (h) Qwen1.5-72B-Chat (API)
+Right Wrong
+response response
+Right tokenization 32 21
+Wrong tokenization 43 135
+Table 30: (i) Chatglm3-6B (Local)
+Right Wrong
+response response
+Right tokenization 59 3
+Wrong tokenization 115 54
+Table 31: (j) Deepseek-R1 (API)
+Appendix C Proportion of four situations between tokenization and response
+Define the four relationships between tokenization and response:
+  • TP: Correct tokenization and correct response.
+  • FP: Incorrect tokenization but correct response.
+  • FN: Correct tokenization but incorrect response.
+  • TN: Incorrect tokenization and incorrect response.
+C.1 ADT-Human (Chinese)
+(a) Baichuan2-13B-Chat (Local)
+(b) Baichuan2-13B-Chat (API)
+(c) Yi-34B-Chat (Local)
+(d) Yi-34B-Chat (API)
+(e) Qwen-7B-Chat (Local)
+(f) Qwen-7B-Chat (API)
+(g) Qwen1.5-72B-Chat (Local)
+(h) Qwen1.5-72B-Chat (API)
+(i) Chatglm3-6B (Local)
+(j) Deepseek-R1 (API)
+C.2 ADT-Human (English)
+(a) Llama-3-8B-Instruct (Local)
+(b) Llama-3-8B-Instruct (API)
+(c) Llama-3-70B-Instruct (Local)
+(d) Llama-3-70B-Instruct (API)
+(e) Mixtral-8x7B-Instruct-v0.1 (Local)
+(f) Mixtral-8x7B-Instruct-v0.1 (API)
+C.3 ADT-Auto
+(a) Baichuan2-13B-Chat (Local)
+(b) Baichuan2-13B-Chat (API)
+(c) Yi-34B-Chat (Local)
+(d) Yi-34B-Chat (API)
+(e) Qwen-7B-Chat (Local)
+(f) Qwen-7B-Chat (API)
+(g) Qwen1.5-72B-Chat (Local)
+(h) Qwen1.5-72B-Chat (API)
+(i) Chatglm3-6B (Local)
+(j) Deepseek-R1 (API)
+Appendix D Details of prompt